OpenStudy (anonymous):

Madeline was completing the square and her work is shown below. Identify the line where she made her mistake. f(x) = 3x^2 + 6x − 7 Line 1: f(x) = 3(x + 2x) − 7 Line 2: f(x) = 3(x + 2x + 1) − 7 − 1 Line 3: f(x) = 3(x + 1)^2 − 8 Line 1 Line 2 Line 3 She did not make any mistakes

OpenStudy (imstuck):

she left off the 3 when she removed the -1 from the parenthesis. Line 2 is where she messed up.

OpenStudy (anonymous):

Thank you! Can you verify for me that this is B? Solve 2x^2 + 12x − 14 = 0 by completing the square. 2(x + 6)^2 = 26; x = −6 + square root 13, x = −6 − square root 13 2(x + 6)^2 = 20; x = −6 + square root 20, x = −6 − square root 20 2(x + 3)^2 = 32; x = 1, x = −7 2(x + 3)^2 = 32; x = 13, x = 19

OpenStudy (imstuck):

yep...just give me a sec

OpenStudy (imstuck):

its the third choice down.

OpenStudy (anonymous):

Thank you

OpenStudy (imstuck):

Do you know how I got that?

OpenStudy (imstuck):

Let me rephrase...do you WANT to know how I got that or do you know already?